s is between r and t, sr=3x+2, st=4x-6, rt=2x+6 I need to find x, sr, st, and rt
Answer by htmentor(1343)

s is between r and t, sr=3x+2, st=4x-6, rt=2x+6 I need to find x, sr, st, and rt
==============================
rs + st has to equal rt.
So we can write 3x + 2 + 4x - 6 = 2x + 6
Solve for x:
7x - 4 = 2x + 6
5x = 10
x = 2
Now substitute the value x = 2 to get the lengths of the line segments.